OpenStudy (anonymous):

in this exercise, find two functions f and g such that (f o g)(x)= h(x). h(x)= (x+3)^2

OpenStudy (anonymous):

what does (f o g) stand for .. is that an OR ?

OpenStudy (anonymous):

its like equal to f(g(x))

OpenStudy (campbell_st):

yes

OpenStudy (campbell_st):

g(x) = x+3 f(x) = x^2 f(g(x)) = (x + 3)^2